# Quasiconvexity in the Riemannian setting

### v3

@@ 1. Introduction
A fundamental result in the calculus of variations concerning the search for minimizers of variational functionals is the characterization of the sequential lower semicontinuity of integral functionals defined on Sobolev spaces $W^{1,p}(\Omega,\mathbb{R}^{m})$, where $\Omega\subset\mathbb{R}^{n}$ is open and bounded. Specifically, for $1\leqslant p\leqslant+\infty$, lower semicontinuity with respect to the weak topology (or weak topology if $p=+\infty$) is equivalent to the quasiconvexity of the integrand, provided certain growth conditions are satisfied.
@@
for every $x\in\mathbb{R}^{n}$, $s\in\mathbb{R}^{m}$ and $\xi\in\mathbb{R}^{{n\times m}}$, where $a:{{\mathbb{R}}}^{n}\to{{\mathbb{R}}}$ is nonnegative and locally summable and $b:{{\mathbb{R}}}^{n}\times\mathbb{R}^{{n\times m}}\to{{\mathbb{R}}}$ is nonnegative and locally bounded.
@@
for every $x\in\mathbb{R}^{n}$, $s\in\mathbb{R}^{m}$ and $\xi\in\mathbb{R}^{{n\times m}}$, where $a:{{\mathbb{R}}}^{n}\to{{\mathbb{R}}}$ is nonnegative and locally summable and $C$ is a nonnegative constant.
@@ 3. Quasiconvexity and semicontinuity
Changing variables as $y=\exp_{x_{0}}^{-1}(x)$, so $y\in r\big(-\frac{1}{2},\frac{1}{2}\big)^{n}\subseteq T_{x_{0}}M$ and denoting with $dy$ the Lebesgue measure on $T_{x_{0}}M$ relative to the metric tensor $g_{x_{0}}$, we have

### v4

@@ 1. Introduction
A fundamental result in the calculus of variations concerning the search for minimizers of variational functionals is the characterization of the sequential lower semicontinuity of integral functionals defined on Sobolev spaces $W^{1,p}(\Omega,\mathbb{R}^{m})$, where $\Omega\subseteq\mathbb{R}^{n}$ is open and bounded. Specifically, for $1\leqslant p\leqslant+\infty$, lower semicontinuity with respect to the weak topology (or weak topology if $p=+\infty$) is equivalent to the quasiconvexity of the integrand, provided certain growth conditions are satisfied.
@@
for almost every $x\in\mathbb{R}^{n}$, $s\in\mathbb{R}^{m}$ and $\xi\in\mathbb{R}^{{n\times m}}$, where $a:{{\mathbb{R}}}^{n}\to{{\mathbb{R}}}$ is nonnegative and locally summable and $b:{{\mathbb{R}}}^{n}\times\mathbb{R}^{{n\times m}}\to{{\mathbb{R}}}$ is nonnegative and locally bounded.
@@
for almost every $x\in\mathbb{R}^{n}$, $s\in\mathbb{R}^{m}$ and $\xi\in\mathbb{R}^{{n\times m}}$, where $a:{{\mathbb{R}}}^{n}\to{{\mathbb{R}}}$ is nonnegative and locally summable and $C$ is a nonnegative constant.
@@ 3. Quasiconvexity and semicontinuity
Changing variables as $y=\exp_{x_{0}}^{-1}(x)$, so $y\in r\big(-\frac{1}{2},\frac{1}{2}\big)^{n}\subseteq T_{x_{0}}M$ and denoting by $dy$ the Lebesgue measure on $T_{x_{0}}M$ relative to the metric tensor $g_{x_{0}}$, we have